OpenStudy (anonymous):

What is the square root of p^6+square root of 9p^6+3 and p^6 square rooted-the square root of p^2...simplify equation should look something like this p^6+9p^6+3p^6-p^2.

OpenStudy (mrdoe):

can you use the equation writer to clean this up my eyes hurt trying to figure out what it says....

OpenStudy (abb0t):

\[p^6 + \sqrt{9p^6+3} \] ??

OpenStudy (anonymous):

\[\sqrt{p^6}+\sqrt{9p^6}+3\sqrt{p^6}-\sqrt{p^2}\]

OpenStudy (anonymous):

the equation looks like that

OpenStudy (abb0t):

First of all, you should know that when you have: \[\sqrt{a^2} = (a^2)^{\frac{ 1 }{ 2} } = a^1 = a\] you can follow a similar process to solve and combine everything and solve.